OpenStudy (anonymous):

how do I get absolute (3x+2)= absolute (10x+3) into standard form?

OpenStudy (anonymous):

\[\left| 3x+2 \right|=\left| 10x+3 \right|\] \[-(3x+2)=\left| 10x+3 \right|\] \[-3x-2=10x+3\]

OpenStudy (tkhunny):

What is "Standard Form"? 10x + 3 > 0 ==> x > -3/10 3x + 2 > 0 ==> x > -2/3 Might be a good place to start.

OpenStudy (anonymous):

sorry I meant to put the standard form where it "=0"